the cojugate of complex number 3+2j is?


1)3-2j


2)2+3j


3)2-3j


4)3+4j

Grade:Upto college level

4 Answers

Neer Varshney
13 years ago

3-2i

vikas askiitian expert
13 years ago

z = 3 + 2i

conjugate of z = 1/z

conjugate z = 1/3+2i

multiplying & dividing by 3-2i

                 = (3-2i/9-4i2)

i2 = -1 so

 conjugate = 3-2i/13

               =3/13 - 2i/13

this is the required ans

approve if u like this ans

vikas askiitian expert
13 years ago

dear vikrant ,

 i have solved for mulltiplicative inverse , that is multiplicative inverse nor conjugate ...

z = x + iy then conjugate is

conjugate = x - iy

real part remains same but sign of  imajinary part is reversed...

Sudheesh Singanamalla
13 years ago

Dear Vikrant ,

to find the conjugate of a complex number a + ib all we need to do is to change the '+' symbol to '-' and if the given complex number is a-ib , we change the '-' to a '+'.

So following that rule

given complex number z = 3 + 2j

so the conjugate of this number is

3 - 2j

option (1) is your answer

 

Please approve